A 100 mol/h liquid oil stream is 20% solute Q. We wish to reduce the concentration of Q to 1% by stripping it out with air. The air flow rate is 100 moles/h. The Henry’s Law constant for Q is 1.5. Find the outlet gas composition and the number of equilibrium stages.
Answer #1

Y2 x = 1 1 11e - 2000 L=100(1-02) = 80 mol/h Y = 1.5% Gs=100 mayno 40 Ls (x2-x)= G CY2-Y) 80(+ - 99) 100(Y2-0) Ya = 0.1919 mo